# Quark-Gluon Plasma, Charmonia, Ultra-relativistic Heavy-Ion Collisions

### v1

@@ Abstract
We provide an update on our semi-classical transport approach for quarkonium production in high-energy heavy-ion collisions, focusing on $J/\psi$ and $\psi(2S)$ mesons in 5.02 TeV Pb-Pb collisions at the Large Hadron Collider (LHC) at both forward and mid-rapidity. In particular, we employ the most recent charm-production cross sections reported in pp collisions, which are pivotal for the magnitude of the regeneration contribution, and their modifications due to cold-nuclear-matter (CNM) effects. Multi-differential observables are calculated in terms of nuclear modification factors as a function of centrality, transverse momentum, and rapidity, including the contributions from bottom-decay feeddown. For our predictions for $\psi(2S)$ production, the mechanism of sequential regeneration relative to the more strongly bound $J/\psi$ meson plays an important role in interpreting recent ALICE data.
@@ 1 Introduction
The production of charmonia in ultra-relativistic heavy-ion collisions (URHICs) has been an active area of research for almost four decades. The initially proposed $J/\psi$ suppression signature of Quark-Gluon Plasma (QGP) formation Matsui and Satz (1986) has developed into more comprehensive transport models that account for regeneration mechanisms as dictated by the principle of detailed balance, see, e.g., Refs. Rapp et al. (2010); Braun-Munzinger and Stachel (2010); Liu et al. (2015); Andronic et al. (2024) for reviews. Abundant charm production at the Large Hadron Collider (LHC), with around 100 charm-anticharm quark pairs in a central Pb-Pb collision at a center-of-mass energy of 5.02 TeV per nucleon pair, has led to predictions of a substantial amount of regenerated charmonia which have been confirmed by experiment. This signature features an approximately constant (or even rising) $J/\psi$ yield with collision centrality in terms of the nuclear modification factor, $R_{\rm AA}$, a concentration of this “extra" yield at relatively low momenta, and an appreciable elliptic flow inherited from charm quarks that are being dragged along with the expanding fireball prior to recombination into charmonium Zhao and Rapp (2011); Song et al. (2011); Zhou et al. (2014); Ferreiro (2015); He et al. (2022). However, significant model uncertainties remain, most notably in the underlying assumptions about the $J/\psi$ dissociation temperature, which controls the onset of regeneration in the cooling fireball, and in the input charm cross section, which determines the equilibrium limit of the charmonia and thus controls the magnitude of the regeneration. The experimental knowledge about the total charm cross section in proton-proton (pp) collisions has much advanced in recent years, with higher precision and an improved assessment of the contribution of charm baryons which has led to a noticeable increase in the value of the cross section Acharya et al. (2022).
@@
Successful measurements of the excited state, $\psi(2S)$, in heavy-ion collisions have been conducted at the Super Proton Synchrotron (SPS) Ramello et al. (1998). Its strong suppression has been explained by both statistical hadronization Andronic et al. (2007), where all charmonium states are produced in relative hadro-chemical equilibrium at the pseudo-critical QCD transition temperature, $T_{\rm pc}\simeq 160$ MeV, and by transport models Sorge et al. (1997); Grandchamp and Rapp (2002). In the latter, the small binding energy of the $\psi(2S)$ (about 60 MeV in vacuum) implies that it has a much smaller dissociation temperature than the $J/\psi$ (with a vacuum binding energy of $\sim$630 MeV), and thus its in-medium kinetics is operative at later stages in the fireball evolution. In this regard, small collision systems, i.e., d-Au(0.2 TeV) collisions at the Relativistic Heavy-Ion Collider (RHIC) Adare et al. (2013) and p-Pb collisions at the LHC Abelev et al. (2014), turned out to give valuable constraints on the reaction rates of the $\psi(2S)$. The expected smaller initial temperatures in these systems did not cause significant $J/\psi$ suppression beyond CNM effects, while the stronger suppression of the $\psi(2S)$ has been interpreted as being due to final-state effects in the more dilute phases of these collisions, relative to Au-Au or Pb-Pb systems Ferreiro (2015). This allowed for a much improved gauge of the $\psi(2S)$ reaction rate Du and Rapp (2015, 2019). When deployed to heavy-ion collisions, this has led to the notion of a “sequential regeneration" of $J/\psi$ and $\psi(2S)$ mesons Du and Rapp (2015). An initial application to CMS data in Pb-Pb (5.02 TeV) collisions at the LHC Khachatryan et al. (2014) involved a cut on the transverse-momentum of the charmonia of $p_{\rm T}\geq 6.5$ GeV, and thus was not directly probing the prevalent regime of the regeneration contributions.
@@
In the present paper, we update our model for quarkonium kinetics in heavy-ion collisions Grandchamp and Rapp (2001); Grandchamp et al. (2004); Zhao and Rapp (2008, 2010) in several respects. Most significantly, we will implement the most recent experimental values for the total charm cross section and refine our treatment of CNM effects (including their $p_{T}$ dependence); we will also utilize an improved input for the in-medium charmonium binding energies as to ensure an approximately constant $J/\psi$ mass, as was recently done in our calculations for bottomonium transport Du et al. (2017), and re-assess the relevance of inelastic-scattering versus gluo-dissociation mechanisms. In our applications to phenomenology, we will specifically elaborate on our previous predictions for recent ALICE data on $\psi(2S)$ production in 5.02 TeV Pb-Pb collisions at the LHC Acharya et al. (2024), thereby reiterating the role that sequential regeneration plays in interpreting these data.
@@
Our paper is organized as follows: In Sec. 2, we briefly recall our calculations of charmonium reaction rates (for $J/\psi$, $\chi_{c}$, and $\psi(2S)$), how they figure in the kinetic-rate equation within a schematic fireball for Pb-Pb collisions at the LHC, and pinpoint updates specific to this work, i.e., the total charm cross section with corrections from nuclear shadowing and bottom-decay feeddown to compute pertinent nuclear modification factors. In Sec. 3 we discuss the time dependence of charmonia kinetics in central and peripheral Pb-Pb collisions at the LHC, as well as the centrality dependence of their inclusive yields with comparisons to data. In Sec. 4, we evaluate the $p_{\rm T}$ dependence of $J/\psi$ and $\psi(2S)$ production in Pb-Pb (5.02 TeV) collisions, based on fits to pp spectra. This analysis encompasses $p_{\rm T}$ spectra, pertinent nuclear modification factors across different centralities, centrality dependent yields within different momentum bins, and the average $p_{\rm T}$ its square, with comparisons to data as available. We summarize and conclude in Sec. 5.
@@ 2.1 Transport Parameters
Our starting point is a kinetic-rate equation that describes the time evolution of charmonium yields, $N_{\psi}$, according to Grandchamp et al. (2004); Zhao and Rapp (2010)
@@
For our purposes below, it is useful to note that the time evolution of $N_{\psi}$ from Eq. 1 can be formally decomposed into two distinct processes corresponding to a primordial and a regeneration component. The primordial component refers to the initial $\psi$ yields that undergo suppression in the medium. This suppression is directly given by the loss term, $-\Gamma_{\psi}(T(\tau))N_{\psi}(\tau)$, reflecting the exponential suppression of the initially produced $\psi$ states. On the other hand, the regeneration component arises from the thermal production of charmonium states within the medium, determined by the approach towards the chemical-equilibrium limit, $N_{\psi}^{\rm eq}(T(\tau))$. The regeneration contribution can thus be defined as the solution to the homogeneous rate equation which starts from a vanishing initial condition. It is equivalent to the difference of the full solution minus the suppression contribution.
@@ 2.2 Charmonium Reaction Rates
The reaction rates in the QGP phase are calculated using in-medium charm-quark masses and binding energies guided by the thermodynamic $T$-matrix approach. In an update to previous work Zhao and Rapp (2010) where results based on the internal-energy potential from lattice-QCD (lQCD) computations were employed Riek and Rapp (2010), we have modified the binding energies to ensure that for the given input charm-quark masses, $m_{c}(T)$, the resulting $J/\psi$ mass is constant with temperature Andronic et al. (2024). This leads to somewhat larger binding energies for temperatures below $\sim$300 MeV, by up to maximum of $\sim$150 MeV at $T$$\simeq$220 MeV, which are, in fact, in better agreement with self-consistent $T$-matrix calculations within the strongly-coupled scenario of Ref. Liu and Rapp (2018) (see, e.g., Ref. Wu et al. (2023)). The net effect on observables, is, however quite small, well within other uncertainties in our in input, such as the nuclear shadowing. The inputs are summarized in Figure 1.
The dominant contribution to the reaction rates arises from inelastic scatterings of thermal partons ($i=q,\bar{q},g)$ with charm quarks inside the bound state, i.e., $i+\psi\to c+\bar{c}+i$. These processes are implemented using perturbative Born diagrams in a quasifree approximation Grandchamp and Rapp (2001), where one of the heavy quarks in the bound state, denoted as $c^{*}$, is assumed to be half-off-shell, thereby carrying the binding energy. The other quark is treated as a spectator, which essentially amounts to neglecting recoil corrections. The dissociation rate is then determined by a convolution of the inelastic (half off-shell) $2\to 2$ cross section (or rather matrix element squared) with a thermal parton distribution function, $f_{i}$,
@@
where $m_{i}$ denotes the thermal-parton mass. The results of the quasifree rates are summarized in Figure 2; they generically show an increase with 3-momentum, mostly caused by a suppression at low momentum, which is quite sensitive to the binding energies, while a weak increase remains even in the limit of vanishing binding due to the increase in final-state phase space. For the $\psi(2S)$, which is essentially unbound even at low QGP temperatures, it turns out Du and Rapp (2015) that coupling the light parton to the bound state using a perturbative scattering diagram is insufficient to describe its suppression observed in d-Au collisions at RHIC Adare et al. (2013) and p-Pb collisions at the LHC Abelev et al. (2014). Therefore, the QGP rates for the $\psi(2S)$ were augmented by a $K$-factor of 3 to simulate nonperturbative interaction strength Du and Rapp (2015).
@@
Figure 2: Dissociation rates of charmonia in the medium as a function of momentum at different temperatures (left panel) and temperature at $p$=0 and 3 GeV(right panel).
We also revisit the relevance of gluo-dissociation processes, $g+\psi\to c+\bar{c}$, whose inelastic rate can be written as Peskin (1979); Bhanot and Peskin (1979),
@@
Here, we adopt a slightly different (and, as we believe, more consistent) implementation compared to our previous studies Zhao and Rapp (2010), by following our recent work on bottomonia Du et al. (2017). Rather than expressing the cross section entirely in terms of its binding energy (as originally derived for a Coulombic bound state), we write it as
@@
respectively. The gluo-dissociation rates of the $J/\psi$ and $\chi_{c}$, shown in Figure 3, are negligible compared to the quasifree rates at temperatures $T\gtrsim 250$ MeV and $T\gtrsim 190$ MeV, respectively (even more so at finite 3-momentum). Once the gluo-dissociation rates become comparable or larger than the quasifree rates at lower temperatures, both rates are numerically already small, implying that their impact on charmonium transport will be small (this will be quantified below). The gluo-dissociation rates for the $\psi(2S)$ (not shown) are negligibly small, as expected from the small binding energy of the $\psi(2S)$.
For the rates in hadronic matter, we follow previous developments Grandchamp and Rapp (2002); Du and Rapp (2015) where effective SU(4) Lagrangian calculations of meson exchange interactions in pion- and $\rho$-meson induced dissociation Lin and Ko (2000); Haglin and Gale (2001) were extended to a large set of non-/strange resonances based on phase space considerations. For the $J/\psi$, the resulting rates remain quite small; however, they are significant for the $\psi(2S)$, although still significantly smaller than in the QGP (with $K$ factor), as shown in the right panel of Figure 2.
@@ 2.3 Charmonium equilibrium limits
where $d_{\psi}$ is the spin degeneracy factor, $E_{\psi}=\sqrt{p^{2}+m_{\psi}^{2}}$ the charmonium energy, $V_{\rm FB}$ denotes the (time-dependent) fireball volume, and $K_{2}$ is the modified Bessel function of the second kind. The charm-quark fugacity, $\gamma_{c}$, is computed as in previous work of our approach Zhao and Rapp (2010); Du et al. (2017), assuming conservation of $c\bar{c}$ pairs during the expansion of the fireball,

### v2

@@ Abstract
We provide an update on our semi-classical transport approach for quarkonium production in high-energy heavy-ion collisions, focusing on $J/\psi$ and $\psi(2S)$ mesons in 5.02 TeV Pb-Pb collisions at the Large Hadron Collider (LHC) at both forward and mid-rapidity. In particular, we employ the most recent charm-production cross sections reported in $p$ collisions, which are pivotal for the magnitude of the regeneration contribution, and their modifications due to cold-nuclear-matter (CNM) effects. Multi-differential observables are calculated in terms of nuclear modification factors as a function of centrality, transverse momentum, and rapidity, including the contributions from feeddown from bottom-hadron decays. For our predictions for $\psi(2S)$ production, the mechanism of sequential regeneration relative to the more strongly bound $J/\psi$ meson plays an important role in interpreting recent ALICE data.
@@ 1 Introduction
The production of charmonia in ultra-relativistic heavy-ion collisions (URHICs) is an active area of research since four decades ago. The initially proposed $J/\psi$ suppression signature of quark-gluon plasma (QGP) formation Matsui:1986dk has developed into more comprehensive transport models that account for regeneration mechanisms as dictated by the principle of detailed balance which ensures that the abundances of charmonia approach their pertinent equilibrium limits, see, e.g., Refs. Rapp:2008tf; Braun-Munzinger:2009dzl; Liu:2015izf; Andronic:2024oxz for reviews. Abundant charm production at the Large Hadron Collider (LHC), with around 100 charm-anticharm quark pairs in a central Pb-Pb collision at a center-of-mass energy of 5.02 TeV per nucleon pair, has led to predictions of a substantial amount of regenerated charmonia which have been confirmed by experiment ALICE:2016flj; Bai:2020svs; ALICE:2023gco; ALICE:2022jeh; ALICE:2019lga; ALICE:2019nrq. This signature features an approximately constant (or even rising) $J/\psi$ yield with collision centrality in terms of the nuclear modification factor, $R_{\rm AA}$, and a concentration of the regeneration yield at relatively low momenta Zhao:2011cv; Song:2011xi; Zhou:2014kka; Ferreiro:2014bia. In addition, the regenerated charmonia exhibit an appreciable elliptic flow inherited from the recombining charm and anti-charm quarks that have been dragged along with the expanding fireball He:2021zej. However, significant model uncertainties remain, most notably in the underlying assumptions about the $J/\psi$ dissociation temperature, which controls the onset of regeneration in the cooling fireball, and in the input charm cross section, which determines the equilibrium limit of the charmonia and thus controls the magnitude of the regeneration. For example, the Statistical Hadronization Model (SHM) offers a complementary perspective, where all charmonium states are produced via statistical hadronization at a fixed common temperature corresponding the pseudo-critical temperature of the chiral cross-over transition, $T_{\rm pc}\simeq 160$ MeV Andronic:2021erx (see also Ref. Capellino:2023cxe), while most transport models are based on a hierarchy of dissociation temperatures that is correlated with the charmonium binding energies. Both transport and SHM models are quite sensitive to the amount of charm-anti-charm quark pairs in the fireball. Fortunately, the experimental knowledge about the total charm cross section in proton-proton ($p$) collisions has much advanced in recent years. With higher precision and an improved assessment of the contribution of charm baryons a noticeable increase in the value of the cross section has emerged ALICE:2021dhb.
@@
Successful measurements of the excited state, $\psi(2S)$, in heavy-ion collisions were conducted at the Super Proton Synchrotron (SPS) RAMELLO1998261c. Its strong suppression has been explained by both the SHM Andronic:2007bi and by transport models Sorge:1997bg; Spieles:1999kp; Grandchamp:2002wp; Linnyk:2008hp. In the latter, the small binding energy of the $\psi(2S)$ (about 60 MeV in vacuum) implies that it has a much smaller dissociation temperature than the $J/\psi$ (with a vacuum binding energy of $\sim$630 MeV), and thus its in-medium kinetics is operative at later stages in the fireball evolution. In this regard, small collision systems, i.e., d-Au(0.2 TeV) collisions at the Relativistic Heavy-Ion Collider (RHIC) PHENIX:2013pmn and p-Pb collisions at the LHC ALICE:2014cgk, turned out to give valuable constraints on the reaction rates of the $\psi(2S)$. The expected smaller initial temperatures in these systems did not cause significant $J/\psi$ suppression beyond CNM effects, while the stronger suppression of the $\psi(2S)$ has been interpreted as being due to final-state effects in the more dilute phases of these collisions, relative to Au-Au or Pb-Pb collision systems Ferreiro:2014bia. This allowed for a much improved gauge of the $\psi(2S)$ reaction rate Du:2015wha; Du:2018wsj. When deployed to heavy-ion collisions, this has led to the notion of a “sequential regeneration" of $J/\psi$ and $\psi(2S)$ mesons Du:2015wha. An initial application to CMS data in Pb-Pb (5.02 TeV) collisions at the LHC CMS:2014vjg involved a cut on the transverse-momentum of the charmonia of $p_{\rm T}\geq 6.5\,{\rm GeV}/c$, and thus was not directly probing the prevalent regime of the regeneration contributions. This has been improved by recent ALICE data ALICE:2022jeh, which will play a key role in what follows below.
@@
In the present paper, we update our model for quarkonium kinetics in heavy-ion collisions Grandchamp:2002wp; Zhao:2010nk in several respects. Most significantly, we will implement the most recent experimental values for the total charm cross section and refine our treatment of CNM effects (including their $p_{T}$ dependence); we will also utilize an improved input for the in-medium charmonium binding energies as to ensure an approximately constant $J/\psi$ mass, as was recently done in our calculations for bottomonium transport Du:2017qkv, and re-assess the relevance of inelastic-scattering versus gluo-dissociation mechanisms. In our applications to phenomenology, we will specifically elaborate on our previous predictions for recent ALICE data on $\psi(2S)$ production in 5.02 TeV Pb-Pb collisions at the LHC ALICE:2022jeh, thereby reiterating the role that sequential regeneration plays in interpreting these data.
@@
This paper is organized as follows: In Sec. 2, we briefly recall our calculations of charmonium reaction rates and how they figure in the kinetic-rate equation within a schematic fireball for Pb-Pb collisions at the LHC; we also pinpoint updates specific to this work, e.g., in-medium charmonium binding energies, the total charm cross section with corrections from nuclear shadowing and bottom-decay feeddown in the nuclear modification factors. In Sec. 3 we discuss the time dependence of charmonia kinetics in central and peripheral Pb-Pb collisions at the LHC, as well as the centrality dependence of their inclusive yields with comparisons to data. In Sec. 4, we evaluate the $p_{\rm T}$ dependence of $J/\psi$ and $\psi(2S)$ production in Pb-Pb (5.02 TeV) collisions, based on fits to $p$ spectra. This analysis encompasses $p_{\rm T}$ spectra, pertinent nuclear modification factors across different centralities, centrality dependent yields within different momentum bins, and the average $p_{\rm T}$ its square, with comparisons to data as available. We summarize and conclude in Sec. 5.
@@ 2.1 Transport Parameters
Our starting point is a kinetic-rate equation that describes the time evolution of charmonium yields, $N_{\psi}$, according to Zhao:2010nk
@@
For our purposes below, it is useful to note that the time evolution of $N_{\psi}$ from Eq. 1 can be formally decomposed into two distinct processes corresponding to a primordial and a regeneration component. The primordial component refers to the initial $\psi$ yields that undergo suppression in the medium. This suppression is directly given by the loss term, $\Gamma_{\psi}(T(\tau))N_{\psi}(\tau)$, reflecting the exponential suppression of the initially produced $\psi$ states. On the other hand, the regeneration component arises from the thermal production of charmonium states within the medium, determined by the approach towards the chemical-equilibrium limit, $N_{\psi}^{\rm eq}(T(\tau))$. The regeneration contribution can thus be defined as the solution to the homogeneous rate equation which starts from a vanishing initial condition. It is equivalent to the difference of the full solution minus the suppression contribution.
@@ 2.2 Charmonium Reaction Rates
The reaction rates in the QGP phase are calculated using in-medium charm-quark masses and binding energies guided by the thermodynamic $T$-matrix approach. In an update to previous work Zhao:2010nk where results based on the internal-energy potential from lattice-QCD (lQCD) computations were employed, we have modified the binding energies to ensure that for the given input charm-quark masses, $m_{c}(T)$, the resulting $J/\psi$ mass is constant with temperature Andronic:2024oxz. This leads to somewhat larger binding energies for temperatures below $\sim$300 MeV, by up to maximum of $\sim$150 MeV at $T\simeq$220 MeV, which are, in fact, in better agreement with self-consistent $T$-matrix calculations within the strongly-coupled scenario of Ref. Liu:2017qah. The net effect on observables, is, however quite small, well within other uncertainties in our in input, such as the nuclear shadowing. The inputs are summarized in Figure 1.
The dominant contribution to the reaction rates arises from inelastic scatterings of thermal partons ($i=q,\bar{q},g)$ with charm quarks inside the bound state, i.e., $i+\psi\to c+\bar{c}+i$. These processes are implemented using perturbative Born diagrams in a quasifree approximation Grandchamp:2002wp, where one of the heavy quarks in the bound state, denoted as $c^{*}$, is assumed to be half-off-shell, thereby carrying the binding energy. The other quark is treated as a spectator, which essentially amounts to neglecting recoil corrections. The dissociation rate is then determined by a convolution of the inelastic (half off-shell) $2\to 2$ cross section (or rather matrix element squared) with a thermal parton distribution function, $f_{i}$,
@@
where $m_{i}$ denotes the thermal-parton mass and $\omega_{c(i)}$ the on-shell energy of the $c$ quark (thermal parton). The results of the quasifree rates are summarized in Figure 2; they generically show an increase with 3-momentum, mostly caused by a suppression at low momentum, which is quite sensitive to the binding energies, while a weak increase remains even in the limit of vanishing binding due to the increase in final-state phase space. For the $\psi(2S)$, which is essentially unbound even at low QGP temperatures, it turns out Du:2015wha that coupling the light parton to the bound state using a perturbative scattering diagram is insufficient to describe its suppression observed in d-Au collisions at RHIC PHENIX:2013pmn and p-Pb collisions at the LHC ALICE:2014cgk. Therefore, the QGP rates for the $\psi(2S)$ were augmented by a $K$-factor of 3 to simulate nonperturbative interaction strength Du:2015wha.
@@
Figure 2: Dissociation rates of charmonia in the medium as a function of momentum at different temperatures (left panel) and temperature at pp=0 and $3\,{\rm GeV}/c$ (right panel).
We also revisit the relevance of gluo-dissociation processes, $g+\psi\to c+\bar{c}$, whose inelastic rate can be written as Peskin:1979va; Bhanot:1979vb,
@@
Here, we adopt a slightly different (and, as we believe, more consistent) implementation compared to our previous studies Zhao:2010nk, by following our recent work on bottomonia Du:2017qkv. Rather than expressing the cross section entirely in terms of its binding energy (as originally derived for a Coulombic bound state), we write it as
@@
respectively. The gluo-dissociation rates of the $J/\psi$ and $\chi_{c}$, shown in Figure 3, are negligible compared to the quasifree rates at temperatures $T\gtrsim 250$ MeV and $T\gtrsim 190$ MeV, respectively (even more so at finite 3-momentum). Once the gluo-dissociation rates become comparable or larger than the quasifree rates at lower temperatures, both rates are numerically small, implying that their impact on charmonium transport will be small (this will be quantified below). The gluo-dissociation rates for the $\psi(2S)$ (not shown) are negligibly small, as expected from the small binding energy of the $\psi(2S)$.
For the rates in hadronic matter, we follow previous developments Grandchamp:2002wp; Du:2015wha where effective SU(4) Lagrangian calculations of meson exchange interactions in pion- and $\rho$-meson induced dissociation Lin:1999ad; Haglin:2000ar were extended to a large set of non-/strange resonances based on phase space considerations. For the $J/\psi$, the resulting rates remain quite small; however, they are significant for the $\psi(2S)$, although still significantly smaller than in the QGP (with $K$ factor), as shown in the right panel of Figure 2.
@@ 2.3 Charmonium equilibrium limits
where $d_{\psi}$ is the spin degeneracy factor, $E_{\psi}=\sqrt{p^{2}+m_{\psi}^{2}}$ the charmonium energy, $V_{\rm FB}$ denotes the (time-dependent) fireball volume, and $K_{2}$ is the modified Bessel function of the second kind. The charm-quark fugacity, $\gamma_{c}$, is computed as in previous work of our approach Zhao:2010nk; Du:2017qkv, assuming conservation of $c\bar{c}$ pairs during the expansion of the fireball,
@@
To account for the non-thermal distributions of charm quarks in the expanding fireballs of URHICs, which tend to suppress the regeneration contribution Du:2022uvj, we adjust the equilibrium limit with a relaxation time factor Grandchamp:2002wp, ${\cal R}=1-\exp\left(-\int\limits_{0}^{\tau}d\tau^{\prime}/\tau_{c}\right)$, where the charm-quark thermalization time, $\tau_{c}$, is taken as $4.5{\rm fm}/c$, representing an approximate average over 3-momentum and temperature (see, e.g., Fig. 3.3 in Ref. He:2022ywp).
2.4 Initial Conditions and Medium Evolution